A sztochasztikus folyamatok előadás hallgatói

dátum:  II. 7. II. 14. II. 21. II. 28. III. 7. III. 14. III. 21.

diákok:

Angyal Orsolya  +
Csikós Diána  +
Csóka Linda  +
Dienes Zoltán  +
Hadházi Ádám  +
Haranghy Szabolcs  +
Juhász Mónika  +
Kása András  +
Kelemen Károly  +
Kontor Ibolya  +
Kovács Tünde  +
Kriczy Tamás  +
Nagy Gergő  +
Nyul Balázs  +
Portik Tamás  +
Vígh Tímea  +